The width of a rectangle is 3 meters less than its length. The area is 28 square meters. Find the length of the rectangle.
Answer by richwmiller(17219)

We know the area of the rectangle so what is the formula for area?
length times width=area
we know the area is 28 square meters.
L*W=28
What do we know about the length and width
We know that the width is the length minus 3
W=L-3
so we substitute L-3 for w
L*(L-3)=28
multiply L times L and L times -3 to get L^2-3L=28
L^2-3L=28
L^2-3L-28=0
How can you solve quadratic equations?
Quadratic formula ,complete the square or by factoring
try factoring
what are the factors of 28
(28,1) (14,2)(7,4) any more?
You want factors of 28 that has one 3 more than the other
Watch out for the signs.
One is + and the other is minus
You can also use the quadratic formula